# Improving the Language Understanding Capabilities of Large Language Models Using Reinforcement Learning

### v2

@@ Abstract
Large language models (LLMs), primarily built on decoder-only transformer architectures, excel in natural language generation tasks and have shown promise in adapting to diverse downstream tasks using zero-shot and few-shot prompting techniques. However, these prompting methods often fall short on natural language understanding (NLU) tasks, where smaller encoder-only models like BERT-base consistently outperform LLMs on benchmarks such as GLUE and SuperGLUE. In this paper, we explore two approaches—supervised fine-tuning and proximal policy optimization (PPO)—to enhance the NLU capabilities of LLMs. To reduce the computational cost of full-model fine-tuning, we integrate low-rank adaptation (LoRA) layers, restricting updates to these layers during both supervised fine-tuning and PPO stages. In the supervised fine-tuning approach, task-specific prompts are concatenated with input queries and ground-truth labels from the NLU training corpus, optimizing the model using the next-token prediction objective. Despite this, LLMs still underperform compared to encoder-only models like BERT-base on several NLU tasks. To address this gap, we employ PPO, a reinforcement learning technique that treats each token generation as an action and evaluates the sequence of generated tokens using a reward function based on their alignment with ground-truth answers. PPO then updates the model to maximize these rewards, effectively aligning its outputs with the correct labels. Our experiments with the LLAMA2-7B model demonstrate that PPO-based fine-tuning significantly improves performance, delivering an average gain of 6.3 points over supervised fine-tuning on the GLUE benchmark. PPO surpasses zero-shot prompting by 38.7 points and few-shot prompting by 26.1 points on GLUE, while also outperforming these baselines by 28.8 and 28.5 points on SuperGLUE. Additionally, PPO exceeds the performance of BERT-large, a strong baseline, with an average improvement of 2.7 points on GLUE and 9.3 points on SuperGLUE. These improvements are consistent across models such as Qwen2.5-7B and MPT-7B, highlighting PPO’s robustness and effectiveness in enhancing the NLU capabilities of LLMs.
@@ 1 Introduction
Large language models (LLMs) Radford et al. (2019); Brown (2020); Touvron et al. (2023b) have revolutionized natural language processing (NLP) with their powerful text generation capabilities, driven by their decoder-only transformer architecture Radford (2018). Pretrained on large amounts of unlabeled text, LLMs can generate coherent and contextually relevant content. Using prompt-based strategies like zero-shot and few-shot prompting Brown (2020), LLMs can tackle various downstream tasks without requiring task-specific fine-tuning. However, these methods often underperform on natural language understanding (NLU) tasks compared to encoder-only models like BERT Devlin (2018), which consistently excel on benchmarks such as GLUE Wang et al. (2019) and SuperGLUE Wang et al. (2020). For instance, our evaluations on LLAMA2-7B showed that zero-shot prompting with task-specific prompts yielded an average performance of 46.1 across all GLUE datasets, while few-shot prompting improved performance to 58.7, both of which significantly lag behind BERT-base’s 79.6 as shown in Table 1. This underperformance is largely due to LLMs’ inability to capture bidirectional context and perform deeper semantic analysis. Enhancing the NLU performance of LLMs remains a challenge, as their autoregressive nature limits their ability to model the bidirectional dependencies crucial for NLU tasks Radford et al. (2019); Brown (2020).
@@
Figure 1: PPO-based fine-tuning of LLAMA2-7B to improve the performance on NLU tasks.
To enhance the performance of LLMs on NLU tasks, we explore two approaches. First, we apply supervised fine-tuning (SFT) of LLMs on NLU training datasets. The model is fine-tuned on input sequences consisting of task-specific prompts, training examples, and their corresponding ground-truth labels, using the next-token prediction objective. To reduce computational overhead, we employ low-rank adaptation (LoRA) layers Hu et al. (2021a), ensuring that only these lightweight matrices are updated during fine-tuning, rather than the entire model. However, in our experiments with LLAMA2-7B, this approach underperforms compared to BERT-base on several GLUE datasets, including QQP, SST-2, STS-B, and MRPC, as detailed in Table 1. On average, across all GLUE datasets, BERT-base achieves a score of 79.6, outperforming LLAMA2-7B fine-tuned model, which attains 78.5. This indicates the need for an alternative approach to further boost performance.
To further enhance the performance of LLMs on NLU tasks, we adopt a proximal policy optimization (PPO) Schulman et al. (2017a) based fine-tuning approach, leveraging LoRA layers to reduce computational complexity. Previous works, including A3C Mnih et al. (2016), AlphaGo Silver et al. (2017b), OpenAI Five OpenAI et al. (2019), and AlphaZero Silver et al. (2017a), have demonstrated that policy-based reinforcement learning can effectively train neural networks to perform actions in complex environments. These methods have also been widely applied to align LLM responses with human preferences Bai et al. (2022a) Ouyang et al. (2022a) and improve reasoning capabilities Havrilla et al. (2024). Building on this foundation, we employ PPO to improve LLM performance on NLU tasks. We frame the task of generating responses by LLMs as a reinforcement learning problem, where the sequence of input tokens represents the state $s_{t}$, and the token generated at each timestep $t$ is treated as the action $a_{t}$. After the entire sequence is generated, a heuristic-based process extracts the answer, which is compared to the ground truth label, and a reward $R$ is assigned accordingly. Our major contributions are:
  • We propose a PPO-based fine-tuning approach to improve the NLU capabilities of LLMs. To reduce computational complexity, we fine-tune only the LoRA layers.
  • Our evaluation on the GLUE and SuperGLUE benchmarks using the LLAMA2-7B model Touvron et al. (2023a) shows that PPO-based fine-tuning significantly outperforms zero-shot and few-shot baselines, with an average improvement of 38.7 and 26.1 points on the GLUE benchmark, and 28.8 and 28.5 points on the SuperGLUE benchmark, respectively. Additionally, PPO-based fine-tuning achieves an average gain of 6.3 points over SFT on GLUE benchmark and outperforms BERT-large, with a 2.7-point gain on GLUE and a 9.3-point improvement on SuperGLUE benchmark.
  • The results are consistent across other LLMs such as Qwen2.5-7B and MPT-7B, demonstrating the robustness of our approach.
@@ 2 Related Works
2.1 Natural Language Understanding
Natural language understanding (NLU) tasks are crucial for evaluating a model’s ability to comprehend and process human language in various contexts, such as classification, inference, and reasoning. The GLUE benchmark Wang et al. (2019) serves as a key standard for NLU performance, covering tasks like CoLA, SST-2, MRPC, MNLI, and so on, which assess grammatical acceptability, sentiment analysis, paraphrase detection, and textual entailment. For more complex challenges, the SuperGLUE benchmark Wang et al. (2020) was introduced, featuring more difficult tasks that require advanced reasoning and comprehension. Together, GLUE and SuperGLUE provide a comprehensive assessment of a model’s language understanding capabilities.
Models such as BERT Devlin (2018), which utilize a bidirectional encoder architecture, have achieved state-of-the-art performance in NLU tasks. BERT’s architecture allows it to capture bidirectional context. Its pretraining strategy, which uses masked language modeling (MLM), helps the model learn deep semantic representations. This combination makes BERT highly effective across a wide range of NLU tasks. The success of encoder-only models in benchmarks such as GLUE and SuperGLUE can largely be attributed to their ability to capture rich bidirectional context during pretraining, which is critical for NLU tasks.
In contrast, LLMs like GPT-2 Radford et al. (2019), GPT-3 Brown (2020), and LLAMA Touvron et al. (2023b) rely on scaling model size with decoder-only architectures, achieving significant success in text generation tasks. However, their zero-shot performance with task-specific prompts remains suboptimal on NLU tasks, such as those in the GLUE benchmark. This underperformance is attributed to their autoregressive nature, which limits their ability to capture the bidirectional dependencies crucial for deep contextual understanding Devlin (2018); Radford et al. (2019); Brown (2020). Efforts to adapt LLMs for NLU have focused on prompt-based methods like few-shot prompting Brown (2020), which show promise but still fall short of the performance achieved by encoder-only models like BERT on these tasks.
2.2 Policy-based Reinforcement Learning
@@
Policy-based reinforcement learning (RL) directly optimizes an agent’s policy by learning its parameters to maximize long-term rewards. Unlike value-based methods like Q-learning Watkins & Dayan (1992) and DQN Hester et al. (2018), which indirectly derive policies through value functions, policy-based methods represent the policy as a parameterized function. This function, $p_{\theta}(a|s)$, defines the probability of taking action $a$ in state $s$, where $\theta$ represents the policy parameters. The goal is to learn optimal parameters $\theta^{*}$ that maximize the expected cumulative reward, typically through policy gradient methods Sutton et al. (1999). These methods excel in high-dimensional or continuous action spaces, where value-based methods can struggle Deisenroth et al. (2013).
@@
Policy-based methods in reinforcement learning (RL) have evolved significantly over time, starting with REINFORCE Williams (1992), which optimizes policies using the policy gradient theorem but suffers from high variance due to its reliance on Monte Carlo estimates of the reward. Monte Carlo estimates refer to calculating the total reward based on full episodes of interaction, meaning updates are made only after an entire sequence of actions and rewards is observed, which can lead to noisy and slow learning. To address this, actor-critic methods like A2C and A3C Mnih (2016) introduced a critic that estimates the value of the current state, allowing for smoother updates by reducing the variability in policy updates and speeding up convergence. However, these methods still faced instability when large updates caused the new policy to diverge too far from the previous one. Trust Region Policy Optimization (TRPO) Schulman (2015) tackled this by limiting the size of policy updates using a KL divergence constraint, but its implementation was complex and computationally expensive. Proximal policy optimization (PPO) Schulman et al. (2017a) simplified this process by introducing a clipped objective function that keeps policy updates within a stable range while being easier to implement. PPO’s balance between simplicity and stability has made it a widely adopted method in modern RL research.
@@
In NLP, PPO has been effectively used in reinforcement learning from human feedback (RLHF) to align LLM outputs with human preferences, as seen in works like InstructGPT Ouyang et al. (2022b) and Constitutional AI Bai et al. (2022b). These approaches treat the LLM as a policy, where model responses are actions, and human feedback serves as rewards. PPO updates the policy based on the reward model trained on human preferences. Additionally, policy-based RL methods have been applied to enhance LLM reasoning capabilities Ziegler et al. (2019); Havrilla et al. (2024); Hu & Shu (2023). In this work, we apply PPO to fine-tune LLMs on NLU tasks.
@@ 3 Preliminaries on Application of PPO for Fine-tuning LLMs
Proximal policy optimization (PPO) Schulman et al. (2017b) is an online reinforcement learning algorithm. In this section, we describe the process to fine-tune an LLM using PPO. During training, at each timestep $t$, the LLM (policy) generates a token prediction $a_{t}$ (action) based on the state $s_{t}$, which consists of the sequence of generated tokens up to timestep $t-1$. The final generated output is evaluated in the context of the downstream task, where the environment provides feedback in the form of rewards. The model updates its parameters based on these rewards to improve its ability to generate accurate predictions over time.
@@
$J(\theta)=\mathbb{E}_{(s_{t},a_{t})\sim\pi_{\theta^{\prime}}}\left[\min\left(\frac{p_{\theta}(a_{t}|s_{t})}{p_{\theta^{\prime}}(a_{t}|s_{t})}\hat{A}_{t},\,\text{clip}\left(\frac{p_{\theta}(a_{t}|s_{t})}{p_{\theta^{\prime}}(a_{t}|s_{t})},\,1-\epsilon,\,1+\epsilon\right)\hat{A}_{t}\right)\right]$
Here, $p_{\theta}(a_{t}|s_{t})$ is the probability of taking action $a_{t}$ in state $s_{t}$ under the current policy, while $p_{\theta^{\prime}}(a_{t}|s_{t})$ represents this probability under the old policy. In PPO, the training data—specifically, the state-action pairs $(s_{t},a_{t})$—are sampled using the old policy $\pi_{\theta^{\prime}}$ (the LLM before it is updated), rather than the new policy currently being optimized. Thus, the ratio $\frac{p_{\theta}(a_{t}|s_{t})}{p_{\theta_{\text{old}}}(a_{t}|s_{t})}$ accounts for how much the new policy has changed relative to the old policy and adjusts the likelihood of an action accordingly. This ratio is multiplied by $\hat{A}_{t}$, the Generalized Advantage Estimation (GAE) Schulman et al. (2018), which measures how much better or worse an action $a_{t}$ is compared to the expected outcome under the current policy.
@@
Here, $R_{t}+\gamma V_{t+1}-V_{t}$ represents the temporal difference (TD) error Sutton (1988). In this expression, $R_{t}$ is the immediate reward received after taking action $a_{t}$, $V_{t}$ is the expected reward before the action, and $\gamma V_{t+1}$ is the discounted estimate of the future reward after the action. This term reflects how the action $a_{t}$ performed when compared to the expected return at state $s_{t}$. The second term, $\gamma\lambda\hat{A}_{t+1}$, is the smoothing factor in GAE, where $\lambda$ is the trade-off parameter. This recursive estimate allows the model to incorporate future information, making the advantage estimate more stable. Smaller values of $\lambda$ emphasize on immediate rewards, while larger values capture longer-term dependencies. The discount factor $\gamma$ controls how much emphasis is placed on future rewards compared to immediate ones, with higher values of $\gamma$ giving more weight to future rewards. $V_{t}$, which represents the expected future reward from state $s_{t}$, is estimated by a critic model.
@@
The critic model consists of a value head, which is a multi-layer perceptron attached to the final layer of the LLM. It takes the LLM’s representation of the generated token sequence up to timestep $t$ (i.e., the state $s_{t}$) and predicts a scalar value representing the value function $V_{t}$ for that state. The critic model is updated using the square of TD error, which is computed as:
@@ 4 Method
To enhance the performance of LLMs on NLU tasks, we adopt two distinct fine-tuning methods. The first approach involves supervised fine-tuning, where the input consists of a concatenation of the task-specific prompt, query and the ground truth answer, with the model optimized using the next-token prediction objective. The second approach utilizes PPO, framing response generation as a reinforcement learning problem. In this setup, the sequence of input tokens till timestep $t-1$ represents the state $s_{t}$, and each token generated at timestep $t$ is treated as an action $a_{t}$. After generating the entire sequence, a heuristic-based process extracts the final answer from this generated sequence, and is compared to the ground truth. PPO is then employed to optimize the model by maximizing the cumulative reward derived from this comparison. To reduce computational complexity, we fine-tune LoRA layers instead of the full model.

### v4

@@ Abstract
Instruction-fine-tuned large language models (LLMs) under 14B parameters continue to underperform on natural language understanding (NLU) tasks, often trailing smaller models like BERT-base on benchmarks such as GLUE and SuperGLUE. Motivated by the success of reinforcement learning in reasoning tasks (e.g., DeepSeek), we explore Proximal Policy Optimization (PPO) as a framework to improve the NLU capabilities of LLMs. We frame NLU as a reinforcement learning environment, treating token generation as a sequence of actions and optimizing for reward signals based on alignment with ground-truth labels. PPO consistently outperforms supervised fine-tuning, yielding an average improvement of 6.3 points on GLUE, and surpasses zero-shot and few-shot prompting by 38.7 and 26.1 points, respectively. Notably, PPO-tuned models outperform GPT-4o by over 4% on average across sentiment and natural language inference tasks, including gains of 7.3% on the Mental Health dataset and 10.9% on SIGA-nli. This work highlights a promising direction for adapting LLMs to new tasks by reframing them as reinforcement learning problems, enabling learning through simple end-task rewards rather than extensive data curation.
Improving the Language Understanding Capabilities of Large Language Models Using Reinforcement Learning
Sai Ashish Somayajulafootnotemark:  Bokai Hu [^1] Equal Contribution  Xin Pan  Pengtao Xie UC San Diego, USA {ssomayaj, p1xie}@ucsd.edu
@@ 1 Introduction
Figure 1: PPO-based fine-tuning of LLAMA2-7B-chat-hf to improve the performance on NLU tasks.
Large language models (LLMs) (Radford et al., 2019; Brown, 2020; Touvron et al., 2023b) have revolutionized natural language processing (NLP) with their powerful text generation capabilities (Radford, 2018). Pretrained on large-scale unlabeled corpora, LLMs can generate coherent and contextually relevant content. Using prompt-based strategies like zero-shot and few-shot prompting (Brown, 2020), these models can address a wide range of downstream tasks without task-specific fine-tuning. However, when applied to instruction-fine-tuned LLMs under 14B parameters—such as LLAMA2-7B-chat-hf—these methods often underperform on natural language understanding (NLU) tasks compared to encoder-only models like BERT (Devlin, 2018), which consistently excel on benchmarks such as GLUE (Wang et al., 2019) and SuperGLUE (Wang et al., 2020). For instance, our evaluation of LLAMA2-7B-chat-hf shows that zero-shot prompting with task-specific prompts yields an average performance of 46.1 across all GLUE datasets, while few-shot prompting improves performance to 58.7—both significantly trailing BERT-base’s 79.6, as shown in Table 1.
To enhance NLU capabilities of LLMs, we investigate reinforcement learning (RL)-based fine-tuning approaches. Motivated by recent work such as DeepSeek Liu et al. (2024), which demonstrates the utility of reward-driven optimization for improving reasoning abilities, we explore the use of Proximal Policy Optimization (PPO) (Schulman et al., 2017a) to align model outputs with task-specific objectives.
While standard fine-tuning (SFT) is commonly used to adapt LLMs to downstream tasks, we find it insufficient for NLU—often underperforming even smaller encoder-only models like BERT-base. In contrast, we use PPO to enhance LLM performance by framing NLU as a reinforcement learning problem. The sequence of input tokens up to timestep $t-1$ represents the state $s_{t}$, and the token generated at timestep $t$ is treated as the action $a_{t}$. After generating the full response, a heuristic extracts the predicted answer, which is compared to the ground-truth label to assign a scalar reward $R$. PPO then updates the model to maximize this reward, enabling direct optimization for task-specific objectives. Empirically, PPO-based fine-tuning of LLAMA2-7B-chat-hf improves GLUE performance by 6.3 points over SFT, surpasses zero- and few-shot prompting by 38.7 and 26.1 points, and even outperforms GPT-4o by over 4% on sentiment and inference tasks—achieving gains of 7.3% on the Mental Health dataset and 10.8% on SIGA-nli. These results demonstrate the effectiveness of reinforcement learning in aligning LLMs under 14B parameters with NLU objectives.
Pre-trained LLMs possess broad linguistic knowledge—spanning syntactic and semantic structures—acquired from large-scale text corpora. We show that reinforcement learning, specifically PPO, can refine this general understanding to better align with task-specific NLU objectives. Similar patterns are observed in DeepSeek, where chain-of-thought pre-training enhances reasoning capabilities, and subsequent RL-based fine-tuning further improves performance. Building on this insight, our findings suggest a promising direction: adapting LLMs to new tasks by formulating them as reinforcement learning problems. When models are sufficiently pre-trained, task alignment may be achieved without additional labeled data—requiring only a well-defined reward function over the outputs. PPO can then optimize the model toward high-reward behaviors. This approach offers a scalable, label-efficient alternative to conventional supervised fine-tuning through reward-driven adaptation.
@@ 2 Related Works
Policy-based reinforcement learning (RL) directly optimizes an agent’s policy by learning its parameters to maximize long-term rewards. Unlike value-based methods like Q-learning (Watkins and Dayan, 1992) and DQN (Hester et al., 2018), which indirectly derive policies through value functions, policy-based methods represent the policy as a parameterized function. This function, $p_{\theta}(a|s)$, defines the probability of taking action $a$ in state $s$, where $\theta$ represents the policy parameters. The goal is to learn optimal parameters $\theta^{*}$ that maximize the expected cumulative reward, typically through policy gradient methods (Sutton et al., 1999). These methods excel in high-dimensional or continuous action spaces, where value-based methods can struggle (Deisenroth et al., 2013).
@@
Policy-based methods in reinforcement learning (RL) have evolved significantly over time, starting with REINFORCE (Williams, 1992), which optimizes policies using the policy gradient theorem but suffers from high variance due to its reliance on Monte Carlo estimates of the reward. Monte Carlo estimates refer to calculating the total reward based on full episodes of interaction, meaning updates are made only after an entire sequence of actions and rewards is observed, which can lead to noisy and slow learning. To address this, actor-critic methods like A2C and A3C (Mnih, 2016) introduced a critic that estimates the value of the current state, allowing for smoother updates by reducing the variability in policy updates and speeding up convergence. However, these methods still faced instability when large updates caused the new policy to diverge too far from the previous one. Trust Region Policy Optimization (TRPO) (Schulman, 2015) tackled this by limiting the size of policy updates using a KL divergence constraint, but its implementation was complex and computationally expensive. Proximal policy optimization (PPO) (Schulman et al., 2017a) simplified this process by introducing a clipped objective function that keeps policy updates within a stable range while being easier to implement. PPO’s balance between simplicity and stability has made it a widely adopted method in modern RL research.
@@
In NLP, PPO has been effectively used in reinforcement learning from human feedback (RLHF) to align LLM outputs with human preferences, as seen in works like InstructGPT (Ouyang et al., 2022) and Constitutional AI (Bai et al., 2022). These approaches treat the LLM as a policy, where model responses are actions, and human feedback serves as rewards. PPO updates the policy based on the reward model trained on human preferences. Additionally, policy-based RL methods have been applied to enhance LLM reasoning capabilities (Ziegler et al., 2019; Havrilla et al., 2024; Hu and Shu, 2023). In this work, we apply PPO to fine-tune LLMs on NLU tasks.
@@ 3 Preliminaries on Application of PPO for Fine-tuning LLMs
Proximal policy optimization (PPO)(Schulman et al., 2017b) is an online reinforcement learning algorithm. In this section, we describe the process to fine-tune an LLM using PPO. During training, at each timestep $t$, the LLM (policy) generates a token prediction $a_{t}$ (action) based on the state $s_{t}$, which consists of the sequence of generated tokens up to timestep $t-1$. The final generated output is evaluated in the context of the downstream task, where the environment provides feedback in the form of rewards. The model updates its parameters based on these rewards to improve its ability to generate accurate predictions over time.
@@
[…]
Here, $p_{\theta}(a_{t}|s_{t})$ is the probability of taking action $a_{t}$ in state $s_{t}$ under the current policy, while $p_{\theta^{\prime}}(a_{t}|s_{t})$ represents this probability under the old policy. In PPO, the training data—specifically, the state-action pairs $(s_{t},a_{t})$—are sampled using the old policy $\pi_{\theta^{\prime}}$ (the LLM before it is updated), rather than the new policy currently being optimized. Thus, the ratio $\frac{p_{\theta}(a_{t}|s_{t})}{p_{\theta_{\text{old}}}(a_{t}|s_{t})}$ accounts for how much the new policy has changed relative to the old policy and adjusts the likelihood of an action accordingly. This ratio is multiplied by $\hat{A}_{t}$, the Generalized Advantage Estimation (GAE) (Schulman et al., 2018), which measures how much better or worse an action $a_{t}$ is compared to the expected outcome under the current policy.
@@
Here, $R_{t}+\gamma V_{t+1}-V_{t}$ represents the temporal difference (TD) error (Sutton, 1988). In this expression, $R_{t}$ is the immediate reward received after taking action $a_{t}$, $V_{t}$ is the expected reward before the action, and $\gamma V_{t+1}$ is the discounted estimate of the future reward after the action. This term reflects how the action $a_{t}$ performed when compared to the expected return at state $s_{t}$. The second term, $\gamma\lambda\hat{A}_{t+1}$, is the smoothing factor in GAE, where $\lambda$ is the trade-off parameter. This recursive estimate allows the model to incorporate future information, making the advantage estimate more stable. Smaller values of $\lambda$ emphasize on immediate rewards, while larger values capture longer-term dependencies. The discount factor $\gamma$ controls how much emphasis is placed on future rewards compared to immediate ones, with higher values of $\gamma$ giving more weight to future rewards. $V_{t}$, which represents the expected future reward from state $s_{t}$, is estimated by a critic model.
@@
The critic model consists of a value head, which is a multi-layer perceptron attached to the final layer of the LLM. It takes the LLMs representation of the generated token sequence up to timestep $t$ (i.e., the state $s_{t}$) and predicts a scalar value representing the value function $V_{t}$ for that state. The critic model is updated using the square of TD error, which is computed as:
@@ 4 Method
To enhance the performance of LLMs on NLU tasks, we adopt two distinct fine-tuning methods. The first approach involves supervised fine-tuning, where the input consists of a concatenation of the task-specific prompt, query and the ground truth answer, with the model optimized using the next-token prediction objective. The second approach utilizes PPO, framing response generation as a reinforcement learning problem. In this setup, the sequence of input tokens until timestep $t-1$ represents the state $s_{t}$, and each token generated at timestep $t$ is treated as an action $a_{t}$. After generating the entire sequence, a heuristic-based process extracts the final answer from this generated sequence, and is compared to the ground truth. PPO is then employed to optimize the model by maximizing the cumulative reward derived from this comparison. To reduce computational complexity, we fine-tune LoRA layers instead of the full model.